OpenStudy (anonymous):

PLease Help!! Does the sequence sin(n) have a convergent subsequence? Please Justify

OpenStudy (anonymous):

Every bounded sequence have a convergent sub-sequence...remember that

OpenStudy (anonymous):

With that you can construct the proof.

OpenStudy (anonymous):

But I'm not sure how to get the subsequence

OpenStudy (anonymous):

well... thats a trivial case lol

OpenStudy (anonymous):

I'm assuming he wants a non-trivial case. but yes that works too

OpenStudy (anonymous):

So it's a rule that sin n = 2*pi*k

OpenStudy (anonymous):

but how did you get n=2*pi*k ?

OpenStudy (anonymous):

dude... look at what 2 Pi k is for sin

OpenStudy (anonymous):

So therefore it oscillates between 1 and -1 as n approaches infinity ?

OpenStudy (anonymous):

And isnt bounded therefore diverges ?

OpenStudy (anonymous):

The Bolzano-Weierstrass Theorem: Every bounded sequence of real numbers has a convergent subsequence.

OpenStudy (anonymous):

I'm assuming n was real

OpenStudy (anonymous):

Ohh... n isn't real... then pellet. lol

OpenStudy (anonymous):

Well try this: Every sequence has a monotone sub-sequence, and if the sequence is bounded, then then monotone sub-sequence is also bounded and converges.

OpenStudy (anonymous):

Alright I get it for this question ... But I'm still confused about the n=2*pi*k .. I dont know what was done with that or how it was gotten.. could you clarify on that please ?

OpenStudy (anonymous):

Due to of course all monotonic bounded sequences converging.... that should be a sufficient proof.

OpenStudy (anonymous):

But it doesnt have a sub sequence in this case correct ?

OpenStudy (anonymous):

The sequence sin(n) is bounded, since -1 ≤ sin(n) ≤ 1 for any value of n. The Bolzano-Weierstrass theorem tells you this must have a convergence subsequence. In general, take any convergent sequence a[n] → L. Then consider any random sequence, probably not convergent, b[n]. Then define the "zipper" of the two to be c[n], where: c[2n] = a[n] c[2n+1] = b[n] Then c[2n] is a convergent subsequence of c[n], since c[2n] = a[n]. Notice how the b[n] could do any sort of whacky thing, and they would not affect this subsequence. So in general, different subsequences can converge to many different things, or diverge, or whatever else -- different kinds of subsequences can "coexist" in a sequence. There are any cases in analysis when a sequence needs to just be well-behaved (e.g. bounded) and then a convergent subsequence must exist. These can be used in other proofs to derive stronger, better theorems.

OpenStudy (anonymous):

Ok I get it .. Thanks guys!!! really appreciate the help !!